# Themis: Towards Flexible and Interpretable NLG Evaluation

### v1

@@ 4.1 Benchmarks
We conducted extensive experiments to evaluate our Themis across six commonly used datasets for different NLG tasks, including SummEval (Fabbri et al., 2021) for summarization, Topical-Chat (Gopalakrishnan et al., 2019) for dialogue response generation, SFRES&SFHOT (Wen et al., 2015) for data-to-text, QAGS (Wang et al., 2020) for factuality, MANS (Guan et al., 2021) for story generation, and WMT23 (Freitag et al., 2023) for machine translation. These datasets were collected into our NLG-Eval corpus but were excluded during the construction of our training data to ensure fairness.
@@ 4.2 Experimental Settings
We chose Llama-3-8B (Meta, 2024) as our base model for supervised fine-tuning and preference alignment, and more training details are described in Appendix C. When assessing the evaluation capability of different models, we calculated the correlation between ratings from model and human using Pearson ($r$), Kendall ($\tau$), and Spearman ($\rho$) correlation coefficients. The specific calculations and the selection of correlation coefficients vary across different datasets, and we follow the common setups from previous work. Furthermore, while previous work often employed a temperature setting of T = 1 to perform multiple sampling and aggregate the evaluation results, which can enhance performance, we believe it is costly in practice and results in the loss of evaluation analyses. Therefore, we tested our Themis with T = 0 to obtain deterministic results, with other methods using their original settings. Given that most metrics and models do not support specific evaluation aspects, we calculate the results of different aspects using their overall scores, as in previous work.
@@ 4.3 Baselines
We experimented with representative and commonly-used evaluation metrics and models from three categories. Traditional metrics include BLEU, ROUGE (Lin, 2004), BARTScore, BERTScore (Zhang et al., 2020), BLEURT (Sellam et al., 2020), CometKiwi (Rei et al., 2022), and UniEval. Methods of direct prompting LLMs include G-Eval and the evaluations with GPT-3.5 and GPT-4, implemented following Chiang and Lee (2023b). Fine-tuned LLMs include X-Eval, Prometheus, Auto-J, TIGERScore, and InstructScore. Many of them are reference-based and therefore cannot be tested on the reference-free QAGS dataset.
@@ 4.4 Main Results
The main experimental results of our Themis compared with other baselines are shown in Table 2. We present the average correlation coefficients for each task, with the complete results demonstrated in Appendix D. Our Themis achieves the best average evaluation performance on six NLG tasks and greatly surpasses other baselines including fine-tuned larger LLMs, except GPT-4. It only outperforms Themis in dialogue response and translation evaluations, likely because GPT-4 has considerably stronger dialogue and multilingual capabilities than Llama-3-8B, which are almost impossible to improve during our training. And our Themis avoids the limitations of high costs and instability in proprietary LLMs and can be conveniently used offline. Furthermore, although models such as BLEURT, X-Eval, and Prometheus conduct evaluation based on references, they still lag behind our reference-free Themis, showing its superiority.
@@ 4.5 Ablation Study
We conducted ablation studies on our two main methods: consistency-based data selection for supervised fine-tuning and rating-guided preference alignment. Additionally, considering the superior performance of Llama 3 among LLMs of the same size, we experimented with other LLMs as the foundational model to verify the effectiveness of our approach. The average results for the six task datasets under different settings are shown in Table 3.
Data Selection We experimented with approximately 67K unfiltered data that has the same scale as the constructed data for supervised fine-tuning, and also trained the model using half and all of the corpus data. The results show that although more data bring some improvements, the training cost increases significantly, and there is still a substantial performance drop compared to our current model.
@@
Foundational Model Although the evaluation models trained based on other LLMs do not perform as well as our current model, the models using Llama-2 still significantly outperformed TIGERScore and Auto-J, which also use Llama-2-13B as the foundation model. It demonstrates the effectiveness of the training data and methods we introduced.
@@ 5 Reliability Analyses
To analyze and investigate the reliability of our Themis, we consider two aspects: generalizability, which means the model can effectively evaluate unseen NLG tasks during training, and robustness, which means the model can accurately evaluate texts with noise and perturbation.
Method Instruction Following Long form Question Answering Average
@@
Table 4: The results on instruction-following and long-form question-answering evaluations.
@@ 5.1 Unseen Tasks
We tested our model on two uncommon and unseen evaluation tasks: long-form question answering and instruction-following evaluation (Li et al., 2023b). The former task involves five evaluation aspects: clarity (CLA), completeness (COM), correctness (COR), politeness (POL), and relevance (REL), while the latter involves example usage (EU), factual accuracy (FA), logical coherence (LC), overall usefulness (OU), and simple language (SL). As shown in Table 4, Themis outperformed both proprietary and open-source LLMs on the overall two tasks, demonstrating good generalization. And its performance is relatively well-balanced without any particular weaknesses, unlike Auto-J, which performs well on instruction-following but poorly on long-form question-answering. Furthermore, some evaluation aspects are relatively open, which traditional reference-based metrics that do not support specific evaluation aspects cannot handle, highlighting the superiority of our Themis.
@@ 5.2 Perturbation Tests
Table 5: The results of perturbation tests.
Hu et al. (2024) has pointed out that both proprietary and open-source LLMs have certain issues with confusing evaluation aspects during NLG evaluation. Therefore, we conducted perturbation tests based on their methods to investigate the robustness of different evaluation models that support customized aspects. Specifically, we applied aspect-targeted perturbations on references in four tasks—news summarization, dialogue summarization, paraphrase generation, and data-to-text—focusing on fluency, coherence, informativeness, and consistency. The perturbations were designed to impact only the quality of the target aspect while leaving other aspects unaffected. We tested the models by comparing ratings on the perturbed texts with those on the original references, expecting little change and decreases in unaffected aspects. As shown in Table 5, Themis is closest to what is expected, with the smallest average decreases across the four tasks. On one hand, slight decreases may be caused by the fact that the designed pertubations do not entirely avoid affecting other aspects. Moreover, the evaluation of perturbed texts may indeed be challenging, necessitating further exploration to improve the evaluator.
@@ 6.1 Prompting LLMs for NLG Evaluation
Previous research has evaluated text quality in various tasks by directly prompting proprietary LLMs (e.g. GPT-4) for scoring Liu et al. (2023b); Chiang and Lee (2023a); Kocmi and Federmann (2023b), comparison (Liusie et al., 2024; Wang et al., 2023d), ranking (Ji et al., 2023; Liu et al., 2023d), and error analysis (Kocmi and Federmann, 2023a; Lu et al., 2023), surpassing traditional evaluation metrics in performance and providing significant flexibility. Leiter et al. (2023) have experimented with various prompt formats and in-contextual example settings. Moreover, to better assess specific aspects, Liu et al. (2024a) have used LLMs to generate or improve the definitions of evaluation criteria for use in prompting, and Gong and Mao (2023a) have had LLMs first evaluate other relevant aspects. Additionally Wu et al. (2023); Bai et al. (2023) have enhanced evaluative capabilities through interactions and role-playing among LLMs. Although these studies are compelling, they rely on proprietary models, which are costly and pose reproducibility issues.
@@ 6.2 Fine-tuned LLM Evaluators
In response to the issues with prompting LLMs, subsequent studies have shifted to fine-tuning relatively smaller open-source large models to create specialized evaluators (Wang et al., 2023e; Xu et al., 2023; Kim et al., 2023; Wang et al., 2023b; Jiang et al., 2023; Li et al., 2023a; Ke et al., 2023; Zhu et al., 2023; Liu et al., 2023a). Their approaches are generally similar: they use existing human evaluation data or synthetic data created by proprietary LLMs, such as GPT-4, as training data to fine-tune open-source LLMs like Llama (Touvron et al., 2023). They vary in details, such as whether references are needed and whether they support customized evaluation criteria. However, without specific measures, such fine-tuned evaluators were shown to contain biases similar to those in proprietary LLMs (Hu et al., 2024).
7 Conclusion
In this paper, we propose a large-scale and comprehensive NLG evaluation corpus and an LLM dedicated to NLG evaluation, Themis. The NLG-Eval corpus encompasses 9 common NLG tasks across 58 datasets, comprising about 0.5 million samples with evaluations from both human and GPT-4. Based on the constructed corpus, we propose a specialized multi-perspective consistency verification method to filter out high-quality supervised fine-tuning data, together with preference alignment guided by evaluation ratings, to train our Themis. It can be applied to various NLG tasks with flexible and interpretable evaluation in a reference-free manner. Extensive experiments demonstrate that Themis performs well on various NLG tasks and can be generalized well to unseen tasks. Our model and resources have been released, and we hope they will promote further research in the field of NLG evaluation.
@@ 7 Conclusion
Due to limited computational resources, we did not use larger and more powerful LLMs than Llama-3-8B as the foundation model for training, such as Llama-3-70B. Although larger models might offer stronger evaluation capabilities, their practical usage becomes inconvenient. Balancing model performance and usability requires further exploration in the future.
@@ Acknowledgements
This work was supported by National Science Foundation of China (No. 62161160339), National Key R&D Program of China (2021YFF0901502), Ant Group Research Fund and Key Laboratory of Science, State Key Laboratory of Media Convergence Production Technology and Systems and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). Xiaojun Wan is the corresponding author.
Appendix A Settings and Prompts for Annotations
We follow the suggestions of Bsharat et al. (2023) to design the instructions and prompts for GPT-4 to annotate NLG evaluation tasks, as shown in Table 18. And similar to Chiang and Lee (2023b), we require to provide the concise and accurate analysis before the evaluation rating with a temperature setting of 1 for sampling. The prompts for re-evaluating evaluations are shown in Table 19 and Table 20, where the task descriptions are appropriately modified to prevent confusion.
Appendix B Details for NLG-Eval Corpus
Our constructed NLG-Eval corpus includes evaluation datasets on nine NLG tasks, including Controllable Generation, Data to Text, Dialogue Response Generation, Grammatical Error Correction, Machine Translation, Paraphrase Generation, Story Generation, Summarization and Text Simplification. The statistics are shown in Table 6 and more detailed meta-information is described in Table 7 to Table 17. Through such data and information, we hope to facilitate related research and advance the field of NLG evaluation.
@@ Appendix B Details for NLG-Eval Corpus
chiang LLM Evaluation (Chiang and Lee, 2023a) 1600 Cohesiveness: How well do the sentences in the story fragment fit together?
@@
Table 7: Datasets on Story Generation Task.
@@
Table 8: Datasets on Controllable Generation Task.
@@
Table 9: Datasets on Data to Text task.
@@
Table 10: Datasets on Dialogue Response Generation task. (Part 1)
@@
Table 11: Datasets on Dialogue Response Generation task. (Part 2)
@@
Table 12: Datasets on Grammatical Error Correction task.
@@
Table 13: Datasets on Paraphrase Generation task.
@@
Table 14: Datasets on Machine Translation task.
@@
Table 15: Datasets on Summarization task. (Part 1)
@@
Table 16: Datasets on Summarization task. (Part 2)
@@
Table 17: Datasets on Text Simplification task.
@@ Appendix C Training Details
The sizes of training data used in our supervised fine-tuning and preference alignment are 67,180 and 10,000, respectively. In the construction of the latter, we prioritize selecting preference pairs with larger differences in evaluation ratings, as intuitively, their preference relations are more reliable. In addition, following the construction method of the training data, we built an additional validation set based on the NLG Eval corpus, comprising a total of 2,479 samples, to find the most suitable training hyperparameters. During supervised fine-tuning, the learning rate is 1e-5, the batch size is 128, the training epoch is 3, and the optimizer is Adam. And in preference alignment, the learning rate is 3e-6, the batch size is 128, the training epoch is 1, $\alpha$ is 1.0, and the optimizer is Adam.
@@ Appendix D Complete Results of Six NLG Tasks
Table 26: Complete results on WMT23 zh-en for machine translation task.

### v2

@@ 4.1 Benchmarks
We conducted extensive experiments to evaluate our Themis across six commonly used evaluation datasets for different NLG tasks, including SummEval (Fabbri et al., 2021) for summarization, Topical-Chat (Gopalakrishnan et al., 2019) for dialogue response generation, SFRES&SFHOT (Wen et al., 2015) for data-to-text, QAGS (Wang et al., 2020) for factuality, MANS (Guan et al., 2021) for story generation, and WMT23 (Freitag et al., 2023) for machine translation (zh-en). These datasets were collected into our NLG-Eval corpus but were excluded during the construction of our training data to ensure fairness.
@@ 4.2 Experimental Settings
We chose Llama-3-8B (Meta, 2024) for supervised fine-tuning and preference alignment in our main experiments, and more training details are described in Appendix C. When assessing the evaluation capability of different models, we calculated the correlation between evaluation ratings from the model and humans using Pearson ($r$), Kendall ($\tau$), and Spearman ($\rho$) correlation coefficients. The specific calculations and the selection of correlation coefficients vary across different datasets, and we follow the common setups used in previous work. Furthermore, while some studies (Chiang and Lee, 2023b) employed a temperature setting of T = 1 to perform multiple sampling and aggregate the evaluation results, which may enhance performance, we believe it is costly and instable in practice and actually leads to the loss of evaluation analyses. Therefore, we tested our Themis with T = 0 and single sampling, while using their original settings for other methods. Given that most metrics and models do not support specific evaluation aspects, we calculate the results of different aspects using their overall scores, as in previous work.
@@ 4.3 Baselines
We experimented with existing representative and commonly-used evaluation metrics and models from three categories. Traditional metrics include BLEU, ROUGE (Lin, 2004), BARTScore, BERTScore (Zhang et al., 2020), BLEURT (Sellam et al., 2020), CometKiwi (Rei et al., 2022), and UniEval. Methods of direct prompting LLMs include G-Eval and the evaluations with GPT-3.5 and GPT-4, implemented following Chiang and Lee (2023b). Fine-tuned evaluation LLMs include X-Eval, Prometheus, Auto-J, TIGERScore, and InstructScore. Many of them are reference-based and therefore cannot be tested on the reference-free QAGS dataset.
@@ 4.4 Main Results
The main experimental results of our Themis compared with other baselines are shown in Table 2. We present the average correlation coefficients for each task, with the complete results demonstrated in Appendix D. Our Themis achieves the best overall evaluation performance and surpasses all larger LLMs. Furthermore, Themis also behaves better on each of six NLG tasks than other baselines, except GPT-4. It only outperforms Themis in dialogue response and translation evaluations, likely because GPT-4 has considerably stronger dialogue and multilingual capabilities than Llama-3-8B, which are almost impossible to improve during our limited training. And our Themis avoids the limitations of high costs and instability in proprietary LLMs and can be conveniently used offline. Moreover, although models such as BLEURT, X-Eval, and Prometheus conduct evaluation based on references, they still lag behind our reference-free Themis, showing its superiority.
@@ 4.5 Ablation Study
We conducted ablation studies on our two main methods: consistency-based data sampling for supervised fine-tuning and rating-guided preference alignment. Additionally, considering the superior performance of Llama-3 among LLMs of the same size, we also experimented with other LLMs as the foundational model to verify the effectiveness of our approach. The average results for the six task datasets with different settings are shown in Table 3.
Data Sampling We experimented with approximately 67K raw data that has the same scale as the constructed data for supervised fine-tuning, and further trained the model using more randomly sampled data, namely half and all of the corpus data. The results show that although more data bring some improvements, the training cost increases significantly, and there is still a substantial performance drop compared to our current model.
@@
Foundational Model Although the evaluation models trained based on other LLMs do not perform as well as our current model, the models fine-tuned on Llama-2-7B still significantly outperformed TIGERScore and Auto-J, which use Llama-2-13B as the foundation model. These results demonstrate the effectiveness and necessity of the data sampling and alignment training methods we introduced.
@@ 5 Reliability Analyses
To analyze and investigate the reliability of our Themis, we consider two aspects: generalizability, which means the model can effectively evaluate unseen NLG tasks during training, and robustness, which means the model can accurately evaluate texts containing noises and perturbations.
Method Instruction Following Long form Question Answering Average $\rho$
@@
Table 4: The spearman correlation between evaluations from humans and different models on instruction-following and long-form question-answering tasks.
@@ 5.1 Unseen Tasks
We tested our model on two common but unseen evaluation tasks during training: long-form question answering and instruction-following evaluation (Li et al., 2023b). The former task involves five evaluation aspects: clarity (CLA), completeness (COM), correctness (COR), politeness (POL), and relevance (REL), while the latter involves example usage (EU), factual accuracy (FA), logical coherence (LC), overall usefulness (OU), and simple language (SL). As shown in Table 4, Themis outperforms both proprietary and open-source LLMs on the overall two tasks, demonstrating good generalization. And its performance is relatively well-balanced without particular weaknesses, unlike Auto-J, which performs well on instruction-following but poorly on long-form question-answering. Furthermore, some evaluation aspects are relatively open, which traditional reference-based metrics that do not support specific evaluation aspects cannot handle, highlighting the superiority of our Themis.
@@ 5.2 Perturbation Tests
Table 5: The results of perturbation tests, showing the changes in evaluation ratings from different models before and after perturbations.
Hu et al. (2024) has pointed out that both proprietary and open-source LLMs have certain issues with confusing evaluation aspects during NLG evaluation. Therefore, we conducted perturbation tests based on their methods to investigate the robustness of different evaluation models that support customized aspects. Specifically, we applied aspect-targeted perturbations on references in four tasks—news summarization, dialogue summarization, paraphrase generation, and data-to-text—focusing on fluency, coherence, informativeness, and consistency. The perturbations were designed to impact only the quality of the target aspect while leaving other aspects unaffected. We tested the models by comparing ratings on the perturbed texts with those on the original references, expecting little change and decreases in unaffected aspects. As shown in Table 5, Themis is closest to what is expected, with the smallest average decreases across the four tasks. And slight decreases may be caused by the fact that the designed pertubations do not entirely avoid affecting other aspects. On the other hand, the evaluation of perturbed texts may indeed be challenging, necessitating further exploration to enhance the model’s capabilities.
@@ 6.1 Prompting LLMs for NLG Evaluation
Previous research has evaluated text quality in various tasks by directly prompting proprietary LLMs (e.g. GPT-4) for scoring Liu et al. (2023b); Chiang and Lee (2023a); Kocmi and Federmann (2023b), comparison (Liusie et al., 2024; Wang et al., 2023d), ranking (Ji et al., 2023; Liu et al., 2023d), and error analysis (Kocmi and Federmann, 2023a; Lu et al., 2023), surpassing traditional evaluation metrics in performance and providing significant flexibility. Leiter et al. (2023) have experimented with various prompt formats and in-contextual example settings. Moreover, to better assess specific aspects, Liu et al. (2024a) have used LLMs to generate or improve the definitions of evaluation criteria for use in prompting, and Gong and Mao (2023a) have had LLMs first evaluate other relevant aspects. Additionally, Wu et al. (2023); Bai et al. (2023) have enhanced evaluative capabilities through interactions and role-playing among LLMs. Although these studies are compelling, they rely on proprietary models, which are costly and pose reproducibility issues.
@@ 6.2 Fine-tuned LLM Evaluators
In response to the issues with prompting LLMs, subsequent studies have shifted to fine-tuning relatively smaller open-source large models to create specialized evaluators (Wang et al., 2023e; Xu et al., 2023; Kim et al., 2023; Wang et al., 2023b; Jiang et al., 2023; Li et al., 2023a; Ke et al., 2023; Zhu et al., 2023; Liu et al., 2023a). Their approaches are generally similar: they use existing human evaluation data or synthetic data created by proprietary LLMs, such as GPT-4, as training data to fine-tune open-source LLMs like Llama (Touvron et al., 2023). Their training modes vary in details, such as whether references are required and whether they support customized evaluation criteria. However, without specific measures, such fine-tuned evaluators seemed to contain issues similar to those in proprietary LLMs (Hu et al., 2024).
7 Conclusions
In this paper, we propose a large-scale and comprehensive NLG evaluation corpus and an LLM dedicated to NLG evaluation, Themis. The NLG-Eval corpus encompasses 9 common NLG tasks across 58 datasets, comprising about 0.5 million samples with evaluations from both human and GPT-4. Based on the constructed corpus, we propose a specialized multi-perspective consistency verification method to select high-quality supervised fine-tuning data, together with preference alignment guided by evaluation ratings, to train our Themis. It can be applied to various NLG tasks with flexible and interpretable evaluation in a reference-free manner. Extensive experiments demonstrate that our Themis performs well on different NLG tasks and can be generalized well to unseen tasks. Our model and resources have been released, and we hope they will promote further research in the field of NLG evaluation.
@@ 7 Conclusions
During the construction of our NLG-Eval corpus, we relied on annotations from GPT-4, resulting in considerable costs. However, it is worthwhile because our Themis, built upon this corpus and proposed training methods, has achieved great NLG evaluation capabilities. We have released the corresponding corpus and model to reduce the cost of future research and promote the development of related studies. Additionally, due to limited computational resources, we did not use larger and more powerful LLMs than Llama-3-8B as the foundational model for training, such as Llama-3-70B. Although larger models might offer stronger evaluation capabilities, their practical usage becomes inconvenient. Balancing model performance and usability requires further exploration in the future.
@@ Acknowledgements
This work was supported by Beijing Science and Technology Program (Z231100007423011), National Science Foundation of China (No. 62161160339), Ant Group Research Fund and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). We appreciate the anonymous reviewers for their helpful comments and all the people who have contributed to this work. Xiaojun Wan is the corresponding author.
Appendix A Details for NLG-Eval Corpus
Our constructed NLG-Eval corpus includes evaluation datasets on nine NLG tasks, including Controllable Generation, Data to Text, Dialogue Response Generation, Grammatical Error Correction, Machine Translation, Paraphrase Generation, Story Generation, Summarization, and Text Simplification. An example from the hsplit dataset for the text simplification task is presented in Figure 3 with simplified illustrations, whose situation is similar for other datasets in NLG-Eval corpus. The statistics are shown in Table 6 and more detailed meta information is described in Table 7 to Table 17. Through such data and information, we hope to facilitate related research and advance the field of NLG evaluation.
@@ Appendix A Details for NLG-Eval Corpus
Chiang-LLM-Evaluation (Chiang and Lee, 2023a) 1600 Cohesiveness: How well do the sentences in the story fragment fit together?
@@
Table 7: Datasets for Story Generation task.
@@
Table 8: Datasets for Controllable Generation task.
@@
Table 9: Datasets for Data to Text task.
@@
Table 10: Datasets for Dialogue Response Generation task. (Part 1)
@@
Table 11: Datasets for Dialogue Response Generation task. (Part 2)
@@
Table 12: Datasets for Grammatical Error Correction task.
@@
Table 13: Datasets for Paraphrase Generation task.
@@
Table 14: Datasets for Machine Translation task.
@@
Table 15: Datasets for Summarization task. (Part 1)
@@
Table 16: Datasets for Summarization task. (Part 2)
@@
Table 17: Datasets for Text Simplification task.
Appendix B Settings and Prompts for GPT-4 Annotations
We follow the suggestions of Bsharat et al. (2023) to design the instructions and prompts for GPT-4 to annotate NLG evaluation tasks, as shown in Table 18. Those for re-evaluating evaluations are shown in Table 19 and Table 20, where the task descriptions are appropriately modified to prevent confusion for GPT-4.
@@ Appendix C Training Details
The sizes of training data used in our supervised fine-tuning and preference alignment are 67,180 and 10,000, respectively. In the construction of the latter, we prioritize selecting preference pairs with larger differences in evaluation ratings, as intuitively, their preference relations are more reliable. In addition, following the construction method of the training data, we built an additional validation set based on the NLG Eval corpus, comprising a total of 2,479 samples, to find the most suitable training hyperparameters. During supervised fine-tuning, the learning rate is 1e-5, the batch size is 128, the training epoch is 3, and the optimizer is AdamW with the zero weight decay. And in preference alignment, the learning rate is 3e-6, the batch size is 128, the training epoch is 1, $\alpha$ is 1.0, and the optimizer is AdamW. Our experiments utilize 8 A100 GPUs during training and inference.
@@ Appendix D Complete Results of Six NLG Tasks
Figure 3: A simplified example from the hsplit dataset for the text simplification task.
@@
Table 26: Complete results on WMT23 (zh-en) for Machine Translation task.